URCES Chapter 02, Problem 018 Z Your answer is par
Answer #1

x = 14t^2 - 5 t^3

v = 28t - 15t^2

a = 28 - 30t

part a

position at t= 6s => x =-576m

as I can see you have solved till part g

so now part h

at the instant particle is not moving ,,,,,so v = 0

we get v=  28t - 15t^2 = 0

t = 28/15 s = 1.8666 s

so at t = 1.8666s we have a = 28 - 30*1.8666 = -28 m/s2

part i

avg vel between t = 0 and t = 6s

avg vel = displacement / time

at t = 0........ x = 0 m

at t = 6s.........x = -576m

so displacement = -576m

avg vel = -576 / 6 = -96m/s